OpenStudy (anonymous):

what is the final concentration of Ammonium nitrate after 250 mL of water are added to a 500 mL of a 0.175 M solution of ammonium nitrate?

OpenStudy (aaronq):

so first write a balanced equation for this, then find how many moles there are of each species.

OpenStudy (anonymous):

the balanced equation for ammonium nitrate is .. HNO3(aq) + NH3(g) -> NH4NO3(aq)

OpenStudy (aaronq):

okay now find the moles of each species present initially. \(Molarity=\dfrac{n_{solute}}{L_{solution}}\)

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

250 mL = 0.25 L and 500 mL = 0.5 L

OpenStudy (aaronq):

find moles

OpenStudy (anonymous):

how do you find moles?

OpenStudy (aaronq):

wait did you change the question? theres no NaClO3 anymore. this is just a dilution: \(M_1V_1=M_2V_2\)

OpenStudy (anonymous):

i messed the question up so i reposted it with the correct info and what do u mean a dilution?

OpenStudy (anonymous):

sorry for the mistake ;/

OpenStudy (aaronq):

it's alright. A dilution means that you're making the solution LESS concentrated. so all you need to do is use the formula above (remember that there is a new volume) \((0.5L)(0.175 M)=(0.25 L + 0.5 L)M_2\) solve for \(M_2\)

OpenStudy (aaronq):

which is the new molarity

OpenStudy (anonymous):

so then it'd be 0.75?

OpenStudy (aaronq):

if thats what you computed then yes

OpenStudy (anonymous):

well i just added 0.25 + 0.5

OpenStudy (aaronq):

yeah thats the new volume, not the new molarity

OpenStudy (anonymous):

so what do i have to do to find the molarity then

OpenStudy (aaronq):

solve for \(M_2\) \((0.5L)(0.175M)=(0.25L+0.5L)M_2\)

OpenStudy (anonymous):

hmm how would i do that

OpenStudy (anonymous):

I'm sorry I've never done a problem like this before

OpenStudy (anonymous):

??

OpenStudy (aaronq):

it's algebra \((0.5L)(0.175M)=(0.25L+0.5L)m_2 \rightarrow M_2=\dfrac{(0.5L)(0.175M)}{(0.25L+0.5L)}\)

OpenStudy (aaronq):

damn they should both say \(M_2\)